# A Comprehensive Survey on Continual Learning in Generative Models

### v1

@@ 4.1.1 Training objectives
4.1.2 Backbone
  • InstructBLIP [213] utilizes a Query Transformer (Q-Former) to extract visual features from a frozen image encoder (ViT-G/14 [209]). The input to the Q-Former contains a set of $K$ learnable query embeddings, which interact with the image encoder’s output through cross attention. The Q-Former outputs are then passed through a linear projection before being fed into the frozen LLM. For the LLM, existing methods [87, 99] typically employ FlanT5-XL which is an instruction-tuned model based on the encoder-decoder Transformer T5 [164]. During training, only the projection layer of the current task in the Q-Former and its corresponding key are trainable.
4.1.3 Benchmark
  • CLiMB [217] is introduced to investigate continual learning in vision-and-language settings, supporting both multimodal and unimodal tasks. It consists of two phases: (1) Upstream Continual Learning of Multimodal Tasks: The pre-trained multimodal model is trained on a sequence of vision-and-language tasks and evaluated after each task to assess Forgetting of previous tasks and Knowledge Transfer to subsequent tasks. (2) Downstream Transfer to Low-Shot Tasks: After each multimodal task, the model is evaluated for its Low-Shot Transfer capability on both multimodal and unimodal downstream tasks. The experiments in CLiMB demonstrates that existing continual learning strategies struggle to generalize across sequences of multimodal tasks and fail to enable effective low-shot adaptation to downstream multimodal or unimodal tasks.
  • VQACL [113] designs dual-level task sequence using two datasets: VQA2.0 [218] and NExT-QA [219]. The outer-level consists of $T$ linguistic-driven tasks, where each task corresponds to learning a specific reasoning skill. For example, for the ’Count’ task, its training data mainly contain the examples that can teach the model how to count, such as ‘How many’ and ‘What number is’. In the inner-level, each linguistic-driven task is further composed of a sequence of visual-driven subtasks, with each subtask containing images from a specific object group. Furthermore, VQACL focuses on evaluating the compositionality of the model’s reasoning skills (e.g., counting, color) and visual concepts (e.g., cat, surfboard). Specifically, a visual-driven subtask is randomly removed from each linguistic-driven task during training and used as a new composition for testing.
  • MLLM-CL [102] introduces two settings named domain continual learning (DCL) and ability continual learning (ACL), providing a comprehensive and realistic evaluation for continual learning of MLLMs. Specifically, DCL focuses on learning domain-specific knowledge continually on several mainstream domains (remote sensing, medical, autonomous driving, science and finance), and evaluated on independently and identically distributed (IID) test sets. Differently, ACL aims to incorporate fundamental abilities (OCR, math & logic, visual perception and GUI agent), evaluating on non-IID scenario.
4.2 Architecture-based Approach
4.2.1 Backbone Expansion
As shown in Figure 5, backbone expansion methods dynamically extend model architectures through modular components, enabling adaptive knowledge acquisition for diverse tasks. Eproj [87] establishes the first continual instruction tuning benchmark for MLLM. To address catastrophic forgetting, the framework dynamically selects optimal Q-Former projection layers through multimodal similarity scoring between input samples and preserved training features, while integrating adaptive regularization constraints with a strategic old-sample replay mechanism to maintain stable model performance. TAM-CL [88] proposes a novel transformer-based continual learning framework that dynamically expands model capacity through task-specific attention layers with learnable tokens. The approach simultaneously preserves acquired knowledge through intermediate teacher-student distillation while employing a compact memory buffer for selective experience replay, ensuring robust performance across sequential tasks.
Figure 5: Illustration of an architecture-based continual learning approach in Multimodal Large Language Models.
@@ 4.2.2 LoRA Expansion
LoRA expansion approaches enhance MLLMs through lightweight low-rank adaptations, where task-specific adapters modules are added to a fixed backbone network, optimally balancing learning plasticity with memory stability for efficient continual learning with minimal forgetting. MoELoRA [89] implements LoRA modules as experts in a mixture-of-experts architecture, employing a learnable router for dynamic expert selection based on task inputs. The approach is evaluated on CoIN, a novel continual instruction tuning benchmark. Continual LLaVA [90] introduces the COAST benchmark, systematically evaluating continual instruction tuning across three dimensions: domain shifts, capability expansion, and dataset evolution. The framework tackles catastrophic forgetting through a novel dual-embedding mechanism, combining instruction-aware intrinsic embeddings with cross-task contextual embeddings, while maintaining parameter efficiency through selective LoRA modules. HiDe-LLaVA [91] proposes a hierarchical decoupling framework for continual instruction tuning in MLLMs. The method analyzes CKA similarity variations across model layers to decompose the architecture into task-specific expansion and task-general fusion, where dual-modality prototype matching guides adaptive LoRA selection while parameter-efficient merging preserves shared knowledge. The work additionally constructs an UCIT benchmark by filtering zero-shot underperforming datasets to eliminate information leakage, enabling more reliable evaluation of catastrophic forgetting mitigation. MR-LoRA [102] employs explicit architectural decoupling for plasticity preservation while preventing catastrophic interference during continual learning. Furthermore, a multimodal routing mechanism is incorporated, which leverages the MLLM’s intrinsic cross-modal representations to dynamically route inputs to specialized parameters, acting as an explicit expert selector in complex multimodal scenarios. LLaCA [92] introduces a continual instruction tuning framework that dynamically balances stability and plasticity via gradient-guided EMA weight adaptation. By leveraging Taylor expansion for optimal weight determination and semantic-aware instruction grouping, the method achieves strong anti-forgetting while preserving new-task performance. SMoLoRA [93] introduces a separable mixture of low-rank adaptation framework to address dual catastrophic forgetting in continual visual instruction tuning. This method employs distinct routing modules for visual understanding and instruction following, dynamically selecting specialized LoRA blocks through adaptive fusion to preserve both capabilities while learning new tasks. Additionally, this work proposes a CVIT benchmark that evaluates model generalization to unseen tasks and robustness under diverse instructions. CL-MoE [94] introduces a Dual-Router MoE that combines instance-level and task-level routing to dynamically select optimal experts, coupled with a Momentum MoE mechanism that employs differential parameter updating strategies for task-shared and task-specific experts to balance knowledge retention and acquisition. LLaVA-CMoE [95] proposes Probe-Guided Knowledge Extension to dynamically expand experts at necessary layers through activation frequency analysis, coupled with a Probabilistic Task Locator that employs hierarchical routing and VAE-based task distribution modeling to prevent interference between new and existing tasks while preserving zero-shot capabilities. FCIT [96] proposes DISCO, a framework for Federated Continual Instruction Tuning in MLLMs, featuring dynamic knowledge organization to manage task-specific subspaces via identity token matching and subspace selective activation to filter irrelevant outputs during inference. It establishes the first FCIT benchmark, addressing real-world challenges of data heterogeneity and continual learning.
@@ 4.2.3 Prompt Expansion
Figure 6: Illustration of an Regularization-based continual learning approach in Multimodal Large Language Models.
ModalPrompt [98] introduces a dual-modality guided prompt learning framework for continual learning of MLLM. It constructs task-specific prototype prompts to retain prior knowledge without data replay and employs a dual-modality (image-text) guided prompt selection mechanism to dynamically match relevant prompts during inference. Fwd-Prompt [99] leveraging prompt tuning with gradient projection to achieve both anti-forgetting and positive forward transfer. It addresses task interference through orthogonal subspace allocation, projecting prompt gradients into residual spaces to preserve old knowledge while reusing pre-trained knowledge in core spaces to enhance future task adaptation. ColPro [100] introduces collaborative prompting, which integrates task-specific question constraint prompting, knowledge acquisition prompting, and visual temporal awareness prompting into a pre-trained LLM. It utilizes a complementary learning principle with General and Expert prompts across LLM layers, employing negative guiding for question awareness, autoregressive temporal dynamics for video understanding, and multimodal distillation to enhance task-specific question answering while mitigating catastrophic forgetting. CluMo [101] proposes a cluster-based modality fusion prompt method, which introduces task-specific visual and textual prompt keys trained via K-means clustering in the first stage and uses their combinations to select optimal prompts from a shared pool during the second stage. It employs a two-stage training strategy with knowledge distillation to enhance generalization and mitigate catastrophic forgetting while maintaining a frozen backbone model.
@@ 4.3.1 Feature-Level Regularization
Another line of work leverages distillation with data from previous tasks or generated pseudo-data. Specifically, MAFED [107] mitigates forgetting by mixing a small portion of old samples into the new task data for distillation. It further introduces a modality-aware feature distillation that dynamically adjusts the distillation weights between question tokens and visual tokens based on their relative importance during training. MoInCL [108] proposes a pseudo targets generation module to synthesize input-target pairs for various task types. These pseudo samples are then mixed with current task data for knowledge distillation, effectively mitigating catastrophic forgetting. QUAD [109] leverages only past task questions for regularization to avoid overfitting to the current task’s answer space. It further introduces attention consistency distillation to preserve both intra-modal and inter-modal attention consistency across tasks.
Figure 7: Illustration of an Replay-based continual learning approach in Multimodal Large Language Models.
@@ 4.3.2 Parameter-Level Regularization
Feature-level distillation strategies preserve prior knowledge by indirectly constraining model updates through additional losses, while parameter-level regularization, as depicted in Figure 6, achieves the same goal by directly regularizing the model parameters. For MLLMs, some methods posit that only a small subset of parameters plays a critical role during downstream task training, while the majority contribute minimally. Consequently, continual learning can be facilitated by selectively protecting these important parameters to mitigate forgetting. For example, Model Tailor [110] first defines two metrics, a sensitivity-based score and a salience-based score, to capture a minimal set of parameters (model patch) that are important when fine-tuning the current downstream task. To enhance target task learning, it further applies a Hessian-based weight compensation method, termed the patch decorator. SPIDER [111] assesses the importance of generalizations in the original weights using weight magnitude analysis and evaluates the importance of specialized knowledge in the fine-tuned model through gradient changes during training. By comparing these two importance distributions, it adaptively fuses the original and fine-tuned weights, effectively preserving the model’s original generic capabilities while incorporating new task knowledge. SEFE [112] categorizes continual learning forgetting in MLLM into superficial forgetting and essential forgetting. To address superficial forgetting, it proposes the answer style diversification paradigm that standardizes response formats across tasks through style transformation. For essential forgetting, it introduces RegLoRA which aggregates positions marked by prior task regularization masks and optimizes them to approach zero, preserving parameters encoding critical prior knowledge.
@@ 4.4 Replay-based Approach
As illustrated in Figure 7, replay techniques in MLLMs generally necessitate the storage of paired image inputs and corresponding instructional data or tokens. For example, VQACL [113] proposes a rehearsal-based method that disentangles sample-specific and sample-invariant features to separate reasoning skills and visual concepts. Sample-specific features are captured via self-attention, while sample-invariant features are learned through dynamically updated prototypes to balance plasticity and stability. SGP [114] leverages scene graphs as concise visual representations to replay pseudo scene graph question answer triplets, where a Symbolic Replay Model based on a language model captures symbolic reasoning mechanisms and learns task-specific mappings between scene graphs and QA pairs, while a Unified VQA Model is trained on both current and replayed samples to adapt to new tasks while retaining prior knowledge, addressing the challenges of multi-modal reasoning and privacy concerns in continual VQA. ProtoGroup [115] proposes a multi-prototype grouping framework, which clusters related prototypes into groups to extract more stable and accurate sample-invariant representations, and employs a memory sample selection method based on prototype group correlations to choose representative and memorable samples for memory updates. Lin et al. [116] leverage memory replay with TF-IDF and K-means clustering to store representative samples, while knowledge distillation guides the model through teacher-student learning to preserve prior knowledge. Additionally, task-specific projection layers are introduced to regularize feature embeddings, minimizing drift across tasks and ensuring continuity in learning, thereby addressing catastrophic forgetting in autonomous driving VQA. Adapt-$\infty$ [117] introduces a dynamic data selection framework that clusters samples into pseudo-skill groups via gradient-based representations, adaptively selects high-impact samples using a multi-way scoring mechanism. To maintain efficiency, it permanently prunes semantically redundant data while preserving task balance. OASIS [118] introduces an adaptive sample selection framework for continual visual instruction tuning. It selects informative samples via forward-pass Fisher Information and reduces redundancy through gradient-based similarity. To ensure efficiency, it dynamically adjusts selection using exponential moving averages without backward computation.
@@ 5 Continual Learning in Vision-Language Action Models
Vision-Language Action (VLA) [220, 221, 20, 222] models refer to exploiting large models to predict action and control behaviors of robots [223, 224] with the aid of large vision language model. It is of great significance to replace the manipulation of humans in certain risky situations and therefore improve the level of automation. However, robots are inherently adapting to a particular field and inevitably encounter the problem of poor generalization [225] when dealing with complex real world. Consequently, continual learning in vision language action [119] aims to equip robots with the ability to adapt and acquire new skills in dynamic and ever-changing real-world scenario like humans do. It is a specific aspect of lifelong robotic learning [226, 227, 228, 229, 230] that employs large multimodal models for robot control and is crucial for robots to accumulate knowledge in novel environments while retaining ability for known skills to mitigate catastrophic forgetting. Figure 8 illustrates the task formulation of visual-language action models within a continual learning framework.

### v2

@@ 4.1.1 Training objectives
4.1.2 Backbone
  • InstructBLIP [201] utilizes a Query Transformer (Q-Former) to extract visual features from a frozen image encoder (ViT-G/14 [202]). The input to the Q-Former contains a set of $K$ learnable query embeddings, which interact with the image encoder’s output through cross-attention. The Q-Former outputs are then passed through a linear projection before being fed into the frozen LLM. For the LLM, existing methods [92, 104] typically employ FlanT5-XL which is an instruction-tuned model based on the encoder-decoder Transformer T5 [168]. During training, only the projection layer of the current task in the Q-Former and its corresponding key are trainable.
Figure 5: Three types of continual learning approaches in Multimodal Large Language Models.
4.1.3 Benchmark
  • CLiMB [206] is introduced to investigate continual learning in vision-and-language settings, supporting both multimodal and unimodal tasks. It consists of two phases: (1) Upstream Continual Learning of Multimodal Tasks: The pre-trained multimodal model is trained on a sequence of vision-and-language tasks and evaluated after each task to assess Forgetting of previous tasks and Knowledge Transfer to subsequent tasks. (2) Downstream Transfer to Low-Shot Tasks: After each multimodal task, the model is evaluated for its Low-Shot Transfer capability on both multimodal and unimodal downstream tasks. The experiments in CLiMB demonstrate that existing continual learning strategies struggle to generalize across sequences of multimodal tasks and fail to enable effective low-shot adaptation to downstream multimodal or unimodal tasks.
  • VQACL [118] designs dual-level task sequence using two datasets: VQA2.0 [207] and NExT-QA [208]. The outer level consists of $T$ linguistic-driven tasks, where each task corresponds to learning a specific reasoning skill. For example, for the ’Count’ task, its training data mainly contain the examples that can teach the model how to count, such as ‘How many’ and ‘What number is’. At the inner level, each linguistic-driven task is further composed of a sequence of visual-driven subtasks, with each subtask containing images from a specific object group. Furthermore, VQACL focuses on evaluating the compositionality of the model’s reasoning skills (e.g., counting, color) and visual concepts (e.g., cat, surfboard). Specifically, a visual-driven subtask is randomly removed from each linguistic-driven task during training and used as a new composition for testing.
  • MLLM-CL [107] introduces two settings named domain continual learning (DCL) and ability continual learning (ACL), providing a comprehensive and realistic evaluation for continual learning of MLLMs. Specifically, DCL focuses on learning domain-specific knowledge continually on several mainstream domains (remote sensing, medical, autonomous driving, science, and finance), and evaluated on independently and identically distributed (IID) test sets. Differently, ACL aims to incorporate fundamental abilities (OCR, math & logic, visual perception and, GUI agent), evaluating non-IID scenarios.
4.2 Architecture-based Approach
4.2.1 Backbone Expansion
As shown in Figure 5 (a), backbone expansion methods dynamically extend model architectures through modular components, enabling adaptive knowledge acquisition for diverse tasks. Eproj [92] establishes the first continual instruction tuning benchmark for MLLM. To address catastrophic forgetting, the framework dynamically selects optimal Q-Former projection layers through multimodal similarity scoring between input samples and preserved training features, while integrating adaptive regularization constraints with a strategic old-sample replay mechanism to maintain stable model performance. TAM-CL [93] proposes a novel transformer-based continual learning framework that dynamically expands model capacity through task-specific attention layers with learnable tokens. The approach simultaneously preserves acquired knowledge through intermediate teacher-student distillation while employing a compact memory buffer for selective experience replay, ensuring robust performance across sequential tasks.
@@ 4.2.2 LoRA Expansion
LoRA expansion approaches enhance MLLMs through lightweight low-rank adaptations, where task-specific adapter modules are added to a fixed backbone network, optimally balancing learning plasticity with memory stability for efficient continual learning with minimal forgetting. MoELoRA [94] implements LoRA modules as experts in a mixture-of-experts architecture, employing a learnable router for dynamic expert selection based on task inputs. The approach is evaluated on CoIN, a novel continual instruction tuning benchmark. Continual LLaVA [95] introduces the COAST benchmark, systematically evaluating continual instruction tuning across three dimensions: domain shifts, capability expansion, and dataset evolution. The framework tackles catastrophic forgetting through a novel dual-embedding mechanism, combining instruction-aware intrinsic embeddings with cross-task contextual embeddings, while maintaining parameter efficiency through selective LoRA modules. HiDe-LLaVA [96] proposes a hierarchical decoupling framework for continual instruction tuning in MLLMs. The method analyzes CKA similarity variations across model layers to decompose the architecture into task-specific expansion and task-general fusion, where dual-modality prototype matching guides adaptive LoRA selection while parameter-efficient merging preserves shared knowledge. The work additionally constructs a UCIT benchmark by filtering zero-shot underperforming datasets to eliminate information leakage, enabling a more reliable evaluation of catastrophic forgetting mitigation. MR-LoRA [107] employs explicit architectural decoupling for plasticity preservation while preventing catastrophic interference during continual learning. Furthermore, a multimodal routing mechanism is incorporated, which leverages the MLLM’s intrinsic cross-modal representations to dynamically route inputs to specialized parameters, acting as an explicit expert selector in complex multimodal scenarios. LLaCA [97] introduces a continual instruction tuning framework that dynamically balances stability and plasticity via gradient-guided EMA weight adaptation. By leveraging Taylor expansion for optimal weight determination and semantic-aware instruction grouping, the method achieves strong anti-forgetting while preserving new-task performance. SMoLoRA [98] introduces a separable mixture of low-rank adaptation framework to address dual catastrophic forgetting in continual visual instruction tuning. This method employs distinct routing modules for visual understanding and instruction following, dynamically selecting specialized LoRA blocks through adaptive fusion to preserve both capabilities while learning new tasks. Additionally, this work proposes a CVIT benchmark that evaluates model generalization to unseen tasks and robustness under diverse instructions. CL-MoE [99] introduces a Dual-Router MoE that combines instance-level and task-level routing to dynamically select optimal experts, coupled with a Momentum MoE mechanism that employs differential parameter updating strategies for task-shared and task-specific experts to balance knowledge retention and acquisition. LLaVA-CMoE [100] proposes Probe-Guided Knowledge Extension to dynamically expand experts at necessary layers through activation frequency analysis, coupled with a Probabilistic Task Locator that employs hierarchical routing and VAE-based task distribution modeling to prevent interference between new and existing tasks while preserving zero-shot capabilities. FCIT [101] proposes DISCO, a framework for Federated Continual Instruction Tuning in MLLMs, featuring dynamic knowledge organization to manage task-specific subspaces via identity token matching and subspace selective activation to filter irrelevant outputs during inference. It establishes the first FCIT benchmark, addressing real-world challenges of data heterogeneity and continual learning.
@@ 4.2.3 Prompt Expansion
ModalPrompt [103] introduces a dual-modality guided prompt learning framework for continual learning of MLLM. It constructs task-specific prototype prompts to retain prior knowledge without data replay and employs a dual-modality (image-text) guided prompt selection mechanism to dynamically match relevant prompts during inference. Fwd-Prompt [104] leverages prompt tuning with gradient projection to achieve both anti-forgetting and positive forward transfer. It addresses task interference through orthogonal subspace allocation, projecting prompt gradients into residual spaces to preserve old knowledge while reusing pre-trained knowledge in core spaces to enhance future task adaptation. ColPro [105] introduces collaborative prompting, which integrates task-specific question constraint prompting, knowledge acquisition prompting, and visual temporal awareness prompting into a pre-trained LLM. It utilizes a complementary learning principle with General and Expert prompts across LLM layers, employing negative guiding for question awareness, autoregressive temporal dynamics for video understanding, and multimodal distillation to enhance task-specific question answering while mitigating catastrophic forgetting. CluMo [106] proposes a cluster-based modality fusion prompt method, which introduces task-specific visual and textual prompt keys trained via K-means clustering in the first stage and uses their combinations to select optimal prompts from a shared pool during the second stage. It employs a two-stage training strategy with knowledge distillation to enhance generalization and mitigate catastrophic forgetting while maintaining a frozen backbone model.
@@ 4.3.1 Feature-Level Regularization
Another line of work leverages distillation with data from previous tasks or generated pseudo-data. Specifically, MAFED [112] mitigates forgetting by mixing a small portion of old samples into the new task data for distillation. It further introduces a modality-aware feature distillation that dynamically adjusts the distillation weights between question tokens and visual tokens based on their relative importance during training. MoInCL [113] proposes a pseudo-target generation module to synthesize input-target pairs for various task types. These pseudo samples are then mixed with current task data for knowledge distillation, effectively mitigating catastrophic forgetting. QUAD [114] leverages only past task questions for regularization to avoid overfitting to the current task’s answer space. It further introduces attention consistency distillation to preserve both intra-modal and inter-modal attention consistency across tasks.
@@ 4.3.2 Parameter-Level Regularization
Feature-level distillation strategies preserve prior knowledge by indirectly constraining model updates through additional losses, while parameter-level regularization, as depicted in Figure 5 (b), achieves the same goal by directly regularizing the model parameters. For MLLMs, some methods posit that only a small subset of parameters plays a critical role during downstream task training, while the majority contribute minimally. Consequently, continual learning can be facilitated by selectively protecting these important parameters to mitigate forgetting. For example, Model Tailor [115] first defines two metrics, a sensitivity-based score and a salience-based score, to capture a minimal set of parameters (model patch) that are important when fine-tuning the current downstream task. To enhance target task learning, it further applies a Hessian-based weight compensation method, termed the patch decorator. SPIDER [116] assesses the importance of generalizations in the original weights using weight magnitude analysis and evaluates the importance of specialized knowledge in the fine-tuned model through gradient changes during training. By comparing these two importance distributions, it adaptively fuses the original and fine-tuned weights, effectively preserving the model’s original generic capabilities while incorporating new task knowledge. SEFE [117] categorizes continual learning forgetting in MLLM into superficial forgetting and essential forgetting. To address superficial forgetting, it proposes the answer style diversification paradigm that standardizes response formats across tasks through style transformation. For essential forgetting, it introduces RegLoRA which aggregates positions marked by prior task regularization masks and optimizes them to approach zero, preserving parameters encoding critical prior knowledge.
@@ 4.4 Replay-based Approach
As illustrated in Figure 5 (c), replay techniques in MLLMs generally necessitate the storage of paired image inputs and corresponding instructional data or tokens. For example, VQACL [118] proposes a rehearsal-based method that disentangles sample-specific and sample-invariant features to separate reasoning skills and visual concepts. Sample-specific features are captured via self-attention, while sample-invariant features are learned through dynamically updated prototypes to balance plasticity and stability. SGP [119] leverages scene graphs as concise visual representations to replay pseudo scene graph question answer triplets, where a Symbolic Replay Model based on a language model captures symbolic reasoning mechanisms and learns task-specific mappings between scene graphs and QA pairs, while a Unified VQA Model is trained on both current and replayed samples to adapt to new tasks while retaining prior knowledge, addressing the challenges of multi-modal reasoning and privacy concerns in continual VQA. ProtoGroup [120] proposes a multi-prototype grouping framework, which clusters related prototypes into groups to extract more stable and accurate sample-invariant representations, and employs a memory sample selection method based on prototype group correlations to choose representative and memorable samples for memory updates. Lin et al. [121] leverage memory replay with TF-IDF and K-means clustering to store representative samples, while knowledge distillation guides the model through teacher-student learning to preserve prior knowledge. Additionally, task-specific projection layers are introduced to regularize feature embeddings, minimizing drift across tasks and ensuring continuity in learning, thereby addressing catastrophic forgetting in autonomous driving VQA. Adapt-$\infty$ [122] introduces a dynamic data selection framework that clusters samples into pseudo-skill groups via gradient-based representations, and adaptively selects high-impact samples using a multi-way scoring mechanism. To maintain efficiency, it permanently prunes semantically redundant data while preserving task balance. OASIS [123] introduces an adaptive sample selection framework for continual visual instruction tuning. It selects informative samples via forward-pass Fisher Information and reduces redundancy through gradient-based similarity. To ensure efficiency, it dynamically adjusts selection using exponential moving averages without backward computation.
@@ 5 Continual Learning in Vision-Language Action Models
Vision-Language Action (VLA) [218, 219, 20, 220] models refer to exploiting large models to predict action and control behaviors of robots [221, 222] with the aid of a large vision language model. It is of great significance to replace the manipulation of humans in certain risky situations and therefore improve the level of automation. However, robots are inherently adapting to a particular field and inevitably encounter the problem of poor generalization [223] when dealing with the complex real world. Consequently, continual learning in vision language action [124] aims to equip robots with the ability to adapt and acquire new skills in dynamic and ever-changing real-world scenarios like humans do. It is a specific aspect of lifelong robotic learning [224, 225, 226, 227, 228] that employs large multimodal models for robot control and is crucial for robots to accumulate knowledge in novel environments while retaining the ability for known skills to mitigate catastrophic forgetting. Figure 6 illustrates the task formulation of visual-language action models within a continual learning framework.